# Arab Spring’s Impact on Science through the Lens of Scholarly Attention, Funding, and Migration

### v2

@@ Acknowledgments
Funding: Yasaman Asgari thanks the University of Zurich and the Digital Society Initiatives for partially funding this project.
@@
Yasaman Asgari, Hongyu Zhou, Özgür Kadir Özer, Rezvaneh Rezapour, Mary Ellen Sloane, and Alexandre Bovet
Corresponding author: Yasaman.asgari@math.uzh.ch
@@
Log Domestic Attention log GDP, log Rpop. Rest of world 0.915 0.042 0.096 0.202
@@
Log Foreign Attention log GDP, log pop. Rest of world 0.933 0.581 0.243 0.436
@@
Log Domestic Funding log GDP Rest of world 0.999 0.158 0.000 0.923
@@
Log Foreign Funding log GDP, log pop. Rest of world 0.125 0.263 0.670 0.005
@@
Log Domestic Funding log GDP Rest of world
@@
Log Foreign Funding log GDP, log pop. Rest of world
@@
Table S5: Difference in differences estimates of the impact of Arab Spring on funding. Estimated coefficients from difference-in-differences regressions evaluating the post-treatment effects of Arab Spring on log funding (total, domestic, foreign). Models control for log GDP and population where applicable. Each colored column represents estimates from separate regressions and compares the target countries with two control groups: the rest of the world and MENA countries less affected by the Arab Spring. Standard errors are in parentheses. ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. $\dagger$ denotes cases where the pre-trend parallel assumption is violated with $p-$value of 5%

### v3

@@ Acknowledgments
Funding: Yasaman Asgari thanks the University of Zurich and the Digital Society Initiative for partially funding this project.
@@
Yasaman Asgari, Hongyu Zhou, Özgür Kadir Özer, Rezvaneh Rezapour, Mary Ellen Sloane, and Alexandre Bovet {Yasaman.asgari, Alexandre.bovet}@uzh.ch
@@
Log Domestic Attention log GDP per capita, log Rpop. Rest of world 0.915 0.042 0.096 0.202
@@
Log Foreign Attention log GDP per capita, log pop. Rest of world 0.933 0.581 0.243 0.436
@@
Log Domestic Funding log GDP per capita Rest of world 0.999 0.158 0.000 0.923
@@
Log Foreign Funding log GDP per capita, log pop. Rest of world 0.125 0.263 0.670 0.005
@@
Log Domestic Funding log GDP per capita Rest of world
@@
Log Foreign Funding log GDP per capita, log pop. Rest of world
@@
Table S5: Difference in differences estimates of the impact of Arab Spring on funding. Estimated coefficients from difference-in-differences regressions evaluating the post-treatment effects of Arab Spring on log funding (total, domestic, foreign). Models control for log GDP and population, where applicable. Each colored column represents estimates from separate regressions and compares the target countries with two control groups: the rest of the world and MENA countries less affected by the Arab Spring. Standard errors are in parentheses. ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. $\dagger$ denotes cases where the pre-trend parallel assumption is violated with $p$-value of 5%
Figure S2: Domestic and foreign scholarly attention before and after the Arab Spring by discipline. The target countries are split into three groups: government overthrow (GO), civil war (CW), and governmental changes (GC). Filled bars represent values before the Arab Spring, while hatched bars indicate values after the Arab Spring. The world’s median before and after the Arab Spring is marked with solid and dashed black lines, respectively. In domestic attention (a–d), social sciences consistently held the smallest values for most countries before the Arab Spring. After the Arab Spring, the social sciences’ scholarly attention on governmental changes has increased. In foreign attention (e–h), social sciences showed notably higher values, particularly in Egypt, Jordan, Syria, and Morocco.
Figure S3: Foreign to Domestic attention ratio before and after the Arab Spring by discipline. Foreign-to-domestic attention ratios are calculated for the four main disciplines, (a) Social Sciences, (b) Health Sciences, (c) Life Sciences, and (d) Physical Sciences, before and after the Arab Spring, for each target country. In Egypt, domestic attention consistently surpassed foreign attention across all disciplines. In countries affected by civil war, the foreign-to-domestic attention ratio increased substantially in the Social Sciences, reaching very high values, and rose from below one to above one in Health Sciences. Meanwhile, the ratio for the governmental changes group displayed varied patterns across disciplines.
Figure S4: Changes in scholarly attention to target countries relative to total scholarly output. The increase in scholarly attention of sources countries strongly correlates with their total scholarly output (a proxy for research capacity and productivity) during the period (Pearson Correlation: 0.684; Spearman Correlation 0.845 both with $p\text{-value}<0.0001$) The USA, Japan, and China have the highest outputs, but the USA shows a significantly greater surge in attention to target countries (172 more papers-fractionally counted-each year). The difference in target scholarly attention is 22 for China and $1$ for Japan.
Figure S5: Shifts in scholarly attention towards target countries as a function of the average GDP per capita. The increase in the contribution of source countries to scholarly attention of target countries correlates with their average GDP per capita between 2002 and 2019 (Pearson Correlation: 0.432; Spearman Correlation 0.608, both with $p\text{-value}<0.0001$, suggesting that economic factors might explain the increased scholarly attention. The contributions of the USA, France, Germany, Italy, Saudi Arabia, and Malaysia to this attention exceeded the global average pattern.
[…]